1. find a equation of the line containing the given pair of points. (-5, -2) and (-3,-1). The equation of the line in slope-intercept form is y=
2. Find an equation of the line containing the given pair of points.(1/4,-1/2) and (3/4,1). y= . Simply. type answer in y=mx+b using integers or fractions.

3. Write an equation of the line containing the given point and parallel to the given line. Express your answer in the form y=mx+b. (8,9); x+4y=9. The equation of the line is y= .

4. Write an equation of the line containing the given point and parallel to the given line. Express your answer in the form y=mx+b. (-7,7); 8x=7y+9.

5. (3,4); 5x+y=7
These are the last ones I need and I just cannot come out with the correct answers. Please someone HELP! THank you.

Answer by Alan3354(69443)

1. find a equation of the line containing the given pair of points. (-5, -2) and (-3,-1).
These are all similar. For this one, find the slope. What did you get for it?
Send your work via the thank you note.
---------------------------

2. Find an equation of the line containing the given pair of points.(1/4,-1/2) and (3/4,1).
Same as #1
-------------------
3. Write an equation of the line containing the given point and parallel to the given line. Express your answer in the form y=mx+b. (8,9); x+4y=9.
Find the slope of the given line first.
The equation of the line is y= . You're wasting time typing that. What purpose does that serve?
-----------------------------

4. Write an equation of the line containing the given point and parallel to the given line. Express your answer in the form y=mx+b. (-7,7); 8x=7y+9.

5. (3,4); 5x+y=7